# HC-Mamba: Vision MAMBA with Hybrid Convolutional Techniques for Medical Image Segmentation

### v3

@@ 1 Introduction
We conducted extensive performance evaluations of the proposed model. The results show that our model has high accuracy (94.84%), mIoU (80.72%) and validity of DSC (89.26%).
@@ 3.2 Results
We compare HC-Mamba with some state-of-the-art models and some recent mamba-based model, presenting the experimental results in Table 1 and 2. In order to fully demonstrate that HC-Mamba’s potential in medical image segmentation tasks directly benefits from SSM, we did not use any pre-training strategies.
For the ISIC2017 and ISIC2018 datasets, HC-Mamba performs well on mIoU and Dice compared to other models. Specifically, HC-Mamba has a 1.59% and 0.9% advantage over MedMamba on mIoU and Dice, respectively, while it has a 2.84% and 1.71% advantage over Unet on mIoU and Dice.
@@
VM-UNet[21] | 77.95 | 87.61 | 94.13 | 96.99 | 85.23 | |
MedMamba[22] | 79.13 | 88.35 | 94.23 | 95.68 | 89.74 | |
HC-Mamba | 80.72 | 89.26 | 94.84 | 97.08 | 88.90 |
For the Synapse datasets, HC-Mamba performs well on DSC and HD95 compared to other models. Specifically, HC-Mamba has a 0.91% and 2.5% advantage over MedMamba and VM-Unet on Dice.
@@
VM-UNet[21] | 79.08 | 32.21 | 84.40 | 67.41 | 84.16 | 80.74 | 92.07 | 56.90 | 87.51 | 79.42 |
MedMamba[22] | 80.67 | 28.15 | 88.23 | 69.53 | 82.22 | 79.49 | 96.89 | 58.57 | 92.29 | 79.12 |
Hc-Mamba | 81.58 | 26.34 | 90.93 | 69.65 | 85.57 | 79.27 | 97.38 | 54.08 | 93.49 | 80.14 |
@@ 3.3 Ablation experiments
- | 77.85 | 87.51 | 27.43 |
Dilated convolution | 78.85 | 88.17 | 24.68 |
DW convolution | 77.95 | 87.61 | 13.06 |
Both | 80.60 | 89.25 | 13.88 |

### v4

@@ 1 Introduction
We conducted extensive performance evaluations of the proposed model. The results show that our model has high accuracy (94.84%), mIoU (78.42%) and validity of DSC (87.89%).
@@ 3.2 Results
We compare HC-Mamba with some state-of-the-art models and some recent mamba-based model, presenting the experimental results in Table 1 and 2.
For the ISIC2017 and ISIC2018 datasets, HC-Mamba performs well on mIoU and Dice compared to other models. Specifically, HC-Mamba has a 0.29% and 0.2% advantage over MedMamba on mIoU and Dice, respectively, while it has a 0.9% and 1.48% advantage over Unet on mIoU and Dice.
@@
| VM-UNet[21] | 77.95 | 87.61 | 94.13 | 96.94 | 85.23 | |
| MedMamba[22] | 78.13 | 87.78 | 94.23 | 95.68 | 89.74 | |
| HC-Mamba | 78.42 | 87.89 | 94.24 | 96.98 | 88.90 |
For the Synapse datasets, HC-Mamba performs well on DSC and HD95 compared to other models. Specifically, HC-Mamba has a 0.31% and 0.5% advantage over MedMamba and VM-Unet on Dice.
@@
| VM-UNet[21] | 79.08 | 32.21 | 84.40 | 68.41 | 83.16 | 80.74 | 92.07 | 56.90 | 87.51 | 79.42 |
| MedMamba[22] | 79.27 | 28.15 | 86.23 | 67.53 | 81.22 | 78.49 | 94.89 | 58.57 | 89.29 | 78.12 |
| Hc-Mamba | 79.58 | 26.34 | 89.93 | 67.65 | 84.57 | 78.27 | 95.38 | 52.08 | 89.49 | 79.84 |
@@ 3.3 Ablation experiments
| - | 77.59 | 87.03 | 27.43 |
| Dilated convolution | 78.40 | 87.85 | 24.68 |
| DW convolution | 77.60 | 87.02 | 13.06 |
| Both | 78.42 | 87.89 | 13.88 |